Which ordered pair is a solution of x+2y=7

Answer by ewatrrr(24785)

x+2y=7
y = -x/2 + 7/2
x = 0 then y = 7/2
(0, 7/2) is one ordered pair that is a solution
that is (0, 7/2) is one of the points on the Line: x + 2y = 7